OpenStudy (rea201):

Prove that g(x)=2x+3/x-1 is the inverse of f(x)=x+3/x-2. I found that it is the inverse just having trouble proving it. I already plugged the function in as x but am having trouble.

hero (hero):

If you're going to write the inverse of a function, you should express the second function as g(x)

hero (hero):

let's let the inverse function be g(x)

hero (hero):

Rewrite g(x) as \[g(x)=1 + \frac{5}{x - 2}\]

OpenStudy (anonymous):

You need to plug the second function as x in the first function, then simplify it; after that, plug the first fuction as x in the second function, then simplify it if both answer are x then you can say that f(x) = g(x) = x therefore they are inverse

hero (hero):

Then find g(f(x))

hero (hero):

as @ktrinh2296 pointed out, if g(x) is truly the inverse of f(x), then f(g(x)) should equal x.

hero (hero):

\[\frac{x + 3}{x - 2} = 1 + \frac{5}{x - 2}\]

hero (hero):

I meant to let that second function equal g(x)

hero (hero):

If we do that then to find g(f(x)), we do this: \[g(f(x)) = 1 + \frac{5}{\frac{2x + 3}{x - 1} -2}\]

OpenStudy (rea201):

how did you get 1+(5/x-2)

hero (hero):

First, before I show you that, I will show you that g(f(x)) = x

OpenStudy (rea201):

ok

hero (hero):

\[g(f(x)) = 1 + 5 \div \left(\frac{2x + 3}{x - 1} - 2\right)\] \[ = 1 + 5 \div \left(\frac{2x + 3 - 2(x - 1)}{x -1}\right)\] \[ = 1 + 5 \div \left(\frac{2x + 3 -2x + 2}{x -1}\right)\] \[ = 1 + 5 \div \left(\frac{2x -2x+ 3 + 2}{x -1}\right)\] \[ = 1 + 5 \div \left(\frac{5}{x -1}\right)\] \[ = 1 + 5 \times \left(\frac{x-1}{5}\right)\] \[ = 1 + \left(\frac{5(x-1)}{5}\right)\] \[ = 1 + \left(\frac{\cancel{5}(x-1)}{\cancel{5}}\right)\] \[ = 1 + (x - 1)\] \[ = 1 + x - 1\] \[ = x + 1 - 1\] \[ = x\]

OpenStudy (rea201):

oh ok

hero (hero):

Now, as promised: \(\large\frac{x + 3}{x - 2} \) \[\\= \frac{x - 2 + 5}{x - 2} \\= \frac{x - 2}{x - 2} + \frac{5}{x - 2}\] \[= 1 + \frac{5}{x-2}\]

OpenStudy (rea201):

ohh ok thank you a bunch. I gave you a metal. If I could have I would have given you 3. :)

OpenStudy (rea201):

I understand it perfect

hero (hero):

You're welcome

hero (hero):

Rewriting g(f(x)) as a whole number plus a fraction makes life much easier when trying to find the composition of functions.

OpenStudy (rea201):

Yes, that is for sure. That is where I made a mistake.

hero (hero):

It wasn't a mistake, you probably just never knew about it. But for reference you should know that something like \[\frac{x + 3}{x -2}\] is actually an improper fraction so it can be re-written as a mixed number.

hero (hero):

If this is homework you have to turn in tomorrow, you'll probably have the best solution in the class.

hero (hero):

But for reference, we let the first function be f(x) and the second function be g(x). As it is, now, you have it backwards.

OpenStudy (rea201):

Just started Calculus. So that should help me. Thanks a lot again.

hero (hero):

It will definitely help you in case you have to integrate a fraction in such form.

OpenStudy (rea201):

Ok again it helped a lot.